# Harnessing Large Language Models for Disaster Management: A Survey

### v1

@@ 3.3.3 Disaster Information Coordination
Coordinating disaster-related information is essential for an organized, collaborative response Comfort et al. (2004); Bharosa et al. (2010). Social media plays a pivotal role, with individuals sharing posts on warnings, needs, and other types of information Lindsay (2011); Imran et al. (2015).
Usefulness Estimation. To enhance the accessibility of useful information, Yamamoto et al. leverage encoder-base LLMs (e.g. BERT and RoBERTa) to filter informative tweets by calculating usefulness ratings Yamamoto et al. (2022).
Relevance Classification. However, it requires a pre-defined threshold to determine tweet relevance. Several studies fine-tune encoder-based LLMs for binary relevance classification, which can also be regarded as disaster identification. Furthermore, BERT is applied to classify posts by relevance with multiple relevance levels Blomeier et al. (2024).
@@
Information Classification. To facilitate information dissemination, several studies fine-tuned encoder-based LLMs to classify posts by information types, including actionable types such as "important for managers" Sharma et al. (2021); humanitarian types such as "Injured people" Yuan et al. (2022); and disaster types Liu et al. (2021).
Several studies introduce advanced techniques to improve performance, with some focusing on augmenting data. Boros et al. augment tweets with hashtags and entities, leveraging RoBERTa for classification and tweet prioritizing Boros et al. (2022). Addressing data label scarcity, Li et al. introduce self-training with BerTweet Nguyen et al. (2020) soft-labeling data for training Li et al. (2021).
Novel frameworks have been introduced to improve performance. For instance, Zou et al. utilize a BERT-Bi-LSTM-CNN model to classify tweets into damage types Zou et al. (2024), with better abilities in capturing temporal dependencies and local patterns. In addition, Zahera et al. utilize a graph attention network (GAT) to capture correlations between tweet embeddings and information types Zahera et al. (2021). Furthermore, FF-BERT combines model strengths with an ensemble of BERT and CNN Wilkho et al. (2024).
Other works enhance LLMs by introducing inductive bias. Nguyen et al. leverage BERTweet to extract rationales (classification supporting evidence) from tweets for humanitarian classification Nguyen and Rudra (2022b) and integrate them with attention mechanisms for better fusion Nguyen and Rudra (2023). RACLC Nguyen and Rudra (2022a) further adopts a two-stage framework, using contrastive learning to refine rationale extraction and facilitate classification.
@@
Furthermore, advanced decoder-based LLMs have been utilized, through instruction-tuning Otal and Canbaz (2024); Yin et al. (2024), zero-shot, and few-shot prompting Dinani et al. (2024) for disaster type and humanitarian classification.
Visual data from social media is further incorporated by multimodal LLMs, where several studies have combined features from BERT and visual models via aggregation Zhang et al. (2022); Yu and Wang (2024). Better fusion strategies have been proposed, where Abavisan et al. propose a cross-attention mechanism to fuse DenseNet image features with BERT text embeddings for humanitarian classification Abavisani et al. (2020). Zhou et al. advance this approach by utilizing dual transformers to fuse ALBERT text features and S-CBAM-VGG image features Zhou et al. (2023a). In addition, Basit et al. classify posts into humanitarian or structural types using the "OR" set of text and image classification results Basit et al. (2023).
@@
Need Classification. Social media also allows individuals to express needs, inspiring LLMs for need identification. Encoder-based LLMs have been utilized to detect disaster needs Yang et al. (2024) and rescue request Toraman et al. (2023). Furthermore, MulTMR employs multiple teacher models with task knowledge to fine-tune BERT and RoBERTa for ranking and detecting actionable help requests Vitiugin and Purohit (2024). Responders also share resources on social media. Lamsal et al. have adopted LLMs to match needs with resources through cosine similarity-based retrieval methods where both offer and request posts are embedded by XLM-RoBERTa Conneau (2019), optimizing resource allocation Lamsal et al. (2024b).
@@
Location Extraction. Additionally, efficient processing of posts enhances information dissemination, especially through location extraction. Several studies fine-tune encoder-based LLMs for location reference recognition (LRR), classifying tokens into categories such as "Inside Locations" (ILOC), and "Other Tokens"(O) Mehmood et al. (2024); Suwaileh et al. (2022). Koshy et al. further improve LRR by adding attention and Bi-LSTM layersKoshy and Elango (2024), while others use a BERT-BiLSTM-CRF model to improve the logic consistency of extracted locations Ma et al. (2022); Zhang et al. (2021). External knowledge bases further aid location extraction. For instance, Caillaut et al. utilize cosine similarity to match post entities with knowledge base Caillaut et al. (2024), improving performance. Decoder-based LLMs are also widely applied, where Yu et al. prompt Llama-3 to extract location-relevant words Yu and Wang (2024). External knowledge has been utilized for more accurate location extraction. Hu et al. leverage GPT-4 for detailed location descriptions, incorporating geo-knowledge within the prompts Hu et al. (2023). In addition, GeoChatGPT augments LLMs with geographic data and Object Character Recognition (OCR)-based object descriptions to predict disaster locations Firmansyah et al. (2024).
Summary Extraction. Furthermore, extracting and summarizing critical posts provides a macro understanding during disasters. Nguyen et al utilize integer linear programming (ILP) with encoder-based LLMs to identify critical posts and rationales for summaries Nguyen and Rudra (2022a); Nguyen et al. (2022). Garg et al. calculate salience scores using BERT embeddings and word2vec key-phrase embeddings from rapid automatic keyword extraction (RAKE)Garg et al. (2024). Decoder-based LLMs go further by generating summaries from retrieved tweets. For instance, Vitiugin et al. rank important tweets utilizing LSTM and then use a T5 model to generate the summaries based on top retrieved tweets Vitiugin and Castillo (2022).
Report Generation. Organizations also use decoder-based LLMs for structured disaster reporting, especially through zero-shot search-based generation Colverd et al. (2023). In addition, Seeberger et al. utilize multiple LLMs each for document retrieval, reranking, and instruction-following summarization Seeberger and Riedhammer (2024b), while chain-of-thought (CoT) reasoning is further utilized Pereira et al. (2023). Building on these methods, Crisis2Sum further organizes query-related information into topic clusters summarized by Mixtral Jiang et al. (2024), and produces a final report by combining key nuggets from each cluster Seeberger and Riedhammer (2024a).
@@ 3.3.4 Disaster Rescuing
Grounded in a comprehensive disaster situation understanding, disaster rescuing aims to save lives and property through timely, coordinated actions.
Plan Generation. Effective rescue operations typically require comprehensive rescue plans. Grounded with rich disaster-related knowledge, decoder-based LLMs have been prompted to generate actionable response plans that provide essential guidance Goecks and Waytowich (2023).
Code Generation. Once a plan is established, decoder-based LLMs can assist in execution performed by organizations and rescue teams, such as assisting in guiding robotic systems by translating verbal inputs into actionable operational commands using RAG Panagopoulos et al. (2024).
@@ 3.3.5 Disaster Issue Consultation
During disasters, affected individuals and organizations often seek reliable guidance. Disaster issue consultation offers advice, safety updates, and expert guidance, thus helping them access resources, evaluate options, and make decisions Jiang (2024).
Answer Generation. Decoder-based LLMs have been prompted to generate answers for frequently asked questions and offer guidance Rawat (2024); Chen and Fang (2024). To mitigate hallucination, LLMs integrate RAG with verified disaster-related documents. For instance, WildfireGPT retrieves wildfire literature and data to augment the prompts Xie et al. (2024). Chen et al. guide LLM reasoning over KG by a prompt chain, adding structured knowledge Chen et al. (2024). Unlike traditional RAG approaches that rely on prompt learning, Xia et al. integrate fine-tuning with RAG to enhance LLMs by combining implicit and external knowledge. Xia et al. (2024). Additionally, multi-modal LLMs can integrate textual and visual data. For instance, several VQA models (e.g., Plug-and-Play VQA) Tiong et al. (2022) have been prompted for zero-shot visual question answering for disaster scenarios Sun et al. (2023). To interpret complex user queries, ADI further introduces sequential modular tools including VLMs, object detection models, and semantic segmentation models Liu et al. (2024). In addition, FloodLense integrates ChatGPT and diffusion models to highlight disaster-affected areas from images to improve flood-related geographical question answering Kumbam and Vejre (2024).
@@ 3.4 Disaster Recovery
LLMs can perform accurate (1) disaster impact assessment, a critical step for disaster recovery. With a comprehensive impact understanding, LLMs can assist decision makers in (2) recovery plans generation. In addition, disaster responders have utilized LLMs for (3) recovery process tracking consistently, ensuring the recovery effectiveness.
@@ 3.4.1 Disaster Impact Assessment
Accurate assessment of the damage degree across physical and social dimensions helps prioritize recovery efforts effectively. Physically, disasters disrupt infrastructure, damage property, and threaten human lives, requiring robust frameworks for the identification and estimation of damage and loss.
Damage Classification. Various encoder-based LLMs have been employed to identify and categorize disaster damage. For instance, Malik et al. utilize BERT to identify the damage-related posts and categorize them as "human damage" or "infrastructure damage" Malik et al. (2024). Chen et al. further categorize damage into four types such as "Water or Power Supply" with BERT Chen and Lim (2021). In addition, Zou et al. propose a BERT-BiLSTM-Sit-CNN framework for damage-related posts identification and damage-type classification with improved textual understanding Zou et al. (2024). Furthermore, several studies utilize LLMs to measure damage severity, where Jeba et al. leverage BERT for damage impact degree classification on social media posts and news  Jeba et al. (2024).
Damage Estimation. Damage severity could be better modeled quantitatively by estimation with a fine-grained understanding. Chen et al. calculate the damage severity scores based on the similarities between post tokens and several seed words’ embeddings, where both embeddings are derived from encoder-based LLMs Chen and Lim (2021).
Answer Generation. In addition, decoder-based LLMs can answer specific assessment questions. For instance, Ziaullah et al. prompt LLMs with RAG to obtain the operation status of critical infrastructure facilities using social media data Ziaullah et al. (2024). Multimodal LLMs further incorporate remote sensing data, with Estevao et al. prompting GPT-4o to generate damage assessment using building images Estêvão (2024). To improve modality alignment, SAM-VQA employs a supervised attention-based VLM to integrate image and question features for VQA tasks Sarkar et al. (2023). Additionally, axillary tasks have been utilized to improve the VQA performance, where DATWEP dynamically balances the importance of segmentation and VQA tasks, adjusting class weights during training Alsan and Arsan (2023).
Statistic Extraction. Furthermore, decoder-based LLMs are prompted to extract fatalities Hou and Xu (2022) from social media via few-shot learning, thus offering timely insights into human loss.
Sentiment Classification. From the social dimension, disasters influence public sentiment and behavior, where encoder-based LLMs Han et al. (2024a); Berbère et al. (2023) have been fine-tuned to classify social media posts into positive and negative emotions. In addition, Li et al. utilize decoder-based LLM (e.g. GPT 3.5) to classify the posts into five emotional types such as "panic" and "sadness" with zero-shot prompting Li et al. (2025), thus aiding responders in understanding and addressing the disaster’s emotional impact.
@@ 3.4.2 Recovery Plan Generation
Based on impact assessment, a recovery plan is developed to rebuild infrastructure, restore services, and enhance resilience Hallegatte et al. (2018).
Plan Generation. Decoder-based LLMs have been utilized in certain recovery scenarios by generating recovery and reconstruction plans. For instance, ChatGPT has been prompted to build disaster recovery plans for business restoration White and Liptak (2024); Lakhera (2024).
@@ 3.4.3 Recovery Process Tracking
Continuous tracking of the recovery process ensures that progress aligns with the planned timeline and goals and decision-makers can adapt recovery strategies to meet evolving needs.
Sentiment Classification. Encoder-based LLM (e.g. BERTweet) has been utilized to assess public sentiment throughout the post-disaster recovery period CONTRERAS et al. , thus enabling responders to tailor their recovery efforts to address emotional needs of affected population effectively.
@@ 4 Datasets
Numerous disaster-related datasets have been utilized to evaluate LLMs in disaster management, and we present a comprehensive list of publicly available datasets in Appendix B.
Classification datasets are primarily designed with textual inputs from platforms like Twitter and news outlets, categorizing data by informativeness (relevance) Olteanu et al. (2014), humanitarian types Imran et al. (2016), damage levels Alam et al. (2021b), and other attributes. Some also include visual data, such as satellite imagery or social media images Alam et al. (2018). Models are evaluated with metrics like accuracy and F1 score.
Estimation datasets include textual and visual inputs and provide quantitative labels such as flood depths Akinboyewa et al. (2024), evaluated using metrics such as Mean Absolute Error (MAE).
Generation datasets are also extensively used, primarily in two categories: question answering and summarization. Question-answering datasets provide disaster-related questions paired with crowd-source annotated answers Rawat (2024). Multimodal question-answering datasets, which incorporate disaster-related images as contextual information, are widely utilized as well Sun et al. (2023). For summarization task, large sets of documents are used as inputs, with reference summaries curated by domain experts McCreadie and Buntain (2023). Both types of generation tasks are evaluated using metrics like ROUGE and BLEU.
Extraction datasets identify and label specific elements within sentence, such as keywords Nguyen and Rudra (2022a) and locations Suwaileh et al. (2022), labeling tokens as “outside,” “start,” or “end” to specify their extraction status. These datasets are primarily used for token-level classification tasks, and is measured with classification metrics.

### v2

@@ 3.3.3 Disaster Information Coordination
Coordinating disaster-related information is crucial for ensuring an organized and collaborative response Comfort et al. (2004); Bharosa et al. (2010). Social media plays a pivotal role in this process, as individuals actively share posts containing warnings, urgent needs, and other critical information Lindsay (2011); Imran et al. (2015).
Usefulness Estimation. To improve the accessibility of valuable information, encoder-base LLMs are utilized to filter informative tweets by computing usefulness ratings Yamamoto et al. (2022). However, this approach requires a predefined threshold to determine the relevance of a tweet.
Relevance Classification. Several studies fine-tune encoder-based LLMs for binary relevance classification, as discussed in Section 3.3.1. Additionally, LLMs have been applied to multi-level relevance classification to further refine disaster-related information filtering Blomeier et al. (2024).
@@
Information Classification. To facilitate information dissemination, several studies have fine-tuned encoder-based LLMs to classify posts based on different information types, including actionable types such as "important for managers" Sharma et al. (2021); humanitarian types such as "Injured people" Yuan et al. (2022); and disaster-specific types Liu et al. (2021). When fine-tuning data is limited, augmentation strategies such as manual hashtag annotation Boros et al. (2022) and self-training with soft labeling Li et al. (2021) are employed to enhance classification performance Lei et al. (2025).
Pure LLM-based methods may have limitations, as discussed in Section 3.3.1. In contrast, hybrid architectures enhance performance by integrating CNNs and BiLSTMs to improve local pattern comprehension Zou et al. (2024) and employing Graph Attention Networks (GATs) to capture correlations between tweet embeddings and information types Zahera et al. (2021). Additionally, FF-BERT leverages an ensemble of BERT and CNN to combine model strengths for improved classification Wilkho et al. (2024). Other studies enhance the application of LLMs in disaster information classification by extracting rationales—evidence that supports classification decisions Nguyen and Rudra (2022b, 2023). RACLC Nguyen and Rudra (2022a) employs a two-stage framework, utilizing contrastive learning to refine rationale extraction and improve classification performance.
@@
(Encoder-)decoder LLMs have also been employed for disaster type and humanitarian classification through instruction tuning Otal and Canbaz (2024); Yin et al. (2024), as well as zero-shot and few-shot prompting Dinani et al. (2024).
Multimodal LLMs can integrate rich data from social media to enhance classification by leveraging multiple modalities Lei et al. (2022). This integration can be achieved through simple feature aggregation Zhang et al. (2022); Yu and Wang (2024) or more advanced fusion techniques, such as cross-attention mechanisms Abavisani et al. (2020) and dual transformer architectures Zhou et al. (2023a). Additionally, Basit et al. classify posts into humanitarian or structural categories only when the text and image classification outputs align; otherwise, the posts are uninformative Basit et al. (2023).
@@
Need Classification. Social media enables individuals to express urgent needs during disasters. Encoder-based LLMs have been employed to detect disaster-related needs Yang et al. (2024); Vitiugin and Purohit (2024) and rescue requests Toraman et al. (2023). Responders also use social media to share available resources. Encoder-based LLMs have been employed to match needs with resources using cosine similarity-based retrieval methods, where both offer and request posts are embedded using XLM-RoBERTa Conneau (2019), optimizing resource allocation.
@@
Location Extraction. Additionally, various post-processing techniques enhance information dissemination, particularly through location extraction. Several studies fine-tune encoder-based LLMs for location reference recognition (LRR), classifying tokens into categories such as "Inside Locations" (ILOC) and "Other Tokens" (O) Mehmood et al. (2024); Suwaileh et al. (2022); Koshy and Elango (2024). LRR can be further improved by integrating a conditional random field (CRF) model, which enhances the logical consistency of extracted locations Ma et al. (2022); Zhang et al. (2021). Furthermore, external knowledge corpora can support location extraction. For instance, Caillaut et al. use cosine similarity to match post entities with a knowledge base, ensuring the authenticity of extracted locations Caillaut et al. (2024).
(Encoder-)decoder LLMs are widely used for extracting location-relevant information through prompt learning Yu and Wang (2024). To enhance accuracy, external knowledge has been incorporated into prompts, including geo-knowledge Hu et al. (2023) and Object Character Recognition-based object descriptions Firmansyah et al. (2024).
Summary Extraction. Furthermore, summarizing disaster-related posts provides a macro-level understanding during crises. Several studies focus on identifying critical and informative posts for summarization by integrating advanced techniques into encoder-based LLMs, such as integer linear programming (ILP) Nguyen and Rudra (2022a); Nguyen et al. (2022) and Rapid Automatic Keyword Extraction (RAKE) Garg et al. (2024).
Summary Generation. (Encoder-)decoder LLMs extend summarization capabilities by generating summaries from retrieved text. For example, Vitiugin et al. rank key tweets using an LSTM model and apply a T5 model to generate summaries based on the top-ranked tweets Vitiugin and Castillo (2022). Crisis2Sum performs query-focused summarization through a multi-step process, including query-informed document retrieval, reranking, fact extraction, clustering, fusion into event nuggets, and final selection for summarization Seeberger and Riedhammer (2024a). Additionally, agent-based approaches can enhance summary quality by leveraging multiple LLMs for document retrieval, reranking, and instruction-following summarization Seeberger and Riedhammer (2024b).
Report Generation. (Encoder-)decoder LLMs have been employed for disaster report generation, utilizing techniques such as RAG to extract relevant web data Colverd et al. (2023) and Chain-of-Thought reasoning to enhance the coherence and accuracy of generated reports.
@@ 3.3.4 Disaster Rescuing
Grounded in a comprehensive understanding of the disaster situation, disaster rescue focuses on saving lives and protecting property through timely and coordinated actions.
Plan Generation. Effective rescue operations require well-structured rescue plans. (Encoder-)decoder LLMs are prompted to generate actionable response plans, offering essential guidance for disaster response Goecks and Waytowich (2023).
Code Generation. Once a plan is established, (encoder-)decoder LLMs can support its execution by assisting organizations and rescue teams. For instance, they can facilitate robotic system guidance during rescue operations by translating verbal inputs into actionable operational commands using RAG Panagopoulos et al. (2024).
@@ 3.3.5 Disaster Issue Consultation
During disasters, affected individuals and organizations often seek reliable guidance. Disaster issue consultation provides advice, safety updates, and expert recommendations, helping them access resources, evaluate options, and make informed decisions Jiang (2024).
Answer Generation. (Encoder-)decoder LLMs are employed to generate answers for frequently asked questions and provide disaster-related guidance Rawat (2024); Chen and Fang (2024). To mitigate hallucination, RAG is integrated with verified disaster-related documents. For example, WildfireGPT retrieves wildfire-related literature and data to enhance prompts Xie et al. (2024). Chen et al. introduce a prompt chain to guide LLM reasoning over a disaster knowledge graph, incorporating structured knowledge Chen et al. (2024). Unlike traditional RAG approaches without training, Xia et al. combine fine-tuning for implicit knowledge updates with RAG for explicit knowledge, further improving response quality Xia et al. (2024).
Additionally, multi-modal LLMs can integrate textual and visual data to enhance disaster response. For example, several visual question answering (VQA) models, such as Plug-and-Play VQA Tiong et al. (2022), have been prompted for zero-shot VQA in disaster scenarios Sun et al. (2023). To handle complex user queries, ADI introduces sequential modular tools, incorporating vision-language models (VLMs), object detection models, and semantic segmentation models Liu et al. (2024). Furthermore, FloodLense combines ChatGPT with diffusion models to highlight disaster-affected areas in images, enhancing flood-related geographical question answering Kumbam and Vejre (2024).
@@ 3.4 Disaster Recovery
LLMs can play a crucial role in (1) disaster impact assessment, a vital step in the recovery process. By providing a comprehensive understanding of disaster impacts, LLMs can assist decision-makers in (2) generating recovery plans tailored to specific needs. Additionally, disaster responders have leveraged LLMs for (3) continuous recovery process tracking, ensuring effectiveness and progress throughout the recovery phase.
@@ 3.4.1 Disaster Impact Assessment
Accurately assessing the extent of damage across both physical and social dimensions is essential for prioritizing recovery efforts effectively.
Damage Classification. From the physical dimension, encoder-based LLMs have been employed to identify and categorize disaster-related damage (e.g., human/infrastructure damage Malik et al. (2024), water/power supply damage Chen and Lim (2021)) Additionally, Zou et al. propose a BERT-BiLSTM-Sit-CNN framework, improving textual understanding for damage-related post identification and damage-type classification Zou et al. (2024). Beyond type classification, LLMs have been utilized to assess damage severity. For instance, Jeba et al. employ BERT to classify damage impact severity in social media posts and news articles Jeba et al. (2024).
Damage Estimation. Damage severity can be more effectively quantified through fine-grained estimation. Chen et al. compute damage severity scores by measuring the similarity between post tokens and predefined seed words’ embeddings, both of which are derived from encoder-based LLMs Chen and Lim (2021).
Answer Generation. In addition, (encoder-)decoder LLMs can answer specific assessment questions. Ziaullah et al. employ RAG-enhanced LLMs to retrieve operational status updates of critical infrastructure facilities from social media data Ziaullah et al. (2024). Multimodal LLMs further incorporate remote sensing data for enhanced assessment. Estevao et al. prompt GPT-4o to generate damage assessments based on building images Estêvão (2024). To improve modality alignment, SAM-VQA employs a supervised attention-based vision-language model (VLM) to integrate image and question features for visual question answering (VQA) tasks Sarkar et al. (2023). Additionally, auxiliary tasks have been leveraged to enhance VQA performance. For instance, DATWEP dynamically balances the significance of segmentation and VQA tasks by adjusting class weights during training Alsan and Arsan (2023).
Statistic Extraction. (Encoder-)decoder LLMs have also used few-shot learning to extract fatality information from social media Hou and Xu (2022), offering timely insights into human loss.
Sentiment Classification. From the social dimension, disasters can influence public sentiment, where encoder-based LLMs Han et al. (2024a); Berbère et al. (2023) have been fine-tuned to classify social media posts into positive and negative emotions. In addition, Li et al. employ (encoder-)decoder LLM (e.g., GPT 3.5) to classify posts into five emotional types, such as "panic" and "sadness", using zero-shot prompting Li et al. (2025). This approach helps responders better understand and address the emotional impact of disasters.
@@ 3.4.2 Recovery Plan Generation
Based on impact assessment, a recovery plan is formulated to rebuild infrastructure, restore services, and strengthen resilience Hallegatte et al. (2018).
Plan Generation. (Encoder-)decoder LLMs have been applied in certain recovery scenarios to generate recovery and reconstruction plans. For example, ChatGPT has been prompted to develop disaster recovery strategies for business restoration White and Liptak (2024); Lakhera (2024).
Figure 2: A sample of dataset for disaster relevance classification from CrisisLexT6 Olteanu et al. (2014).
@@ 3.4.3 Recovery Process Tracking
Continuous tracking of the recovery process ensures that progress remains aligned with the planned timeline, allowing decision-makers to adapt recovery strategies to evolving needs.
Sentiment Classification. Encoder-based LLMs (e.g., BERTweet) have been employed to assess public sentiment throughout the recovery period CONTRERAS et al. , enabling responders to tailor recovery efforts to effectively address the emotional needs of affected populations.
Figure 3: Publication counts utilizing existing LLMs and developing new models across the four phases of disaster management.
@@ 4 Datasets
Multiple disaster-related datasets have been employed to evaluate LLMs in disaster management. A comprehensive list of publicly available datasets is provided in Appendix B.
Classification datasets primarily consist of textual inputs from platforms such as Twitter and news outlets, categorizing data based on informativeness (relevance) Olteanu et al. (2014) (illustrated in Figure 2), humanitarian types Imran et al. (2016), damage levels Alam et al. (2021b), and other relevant attributes. Some datasets also incorporate visual data, including satellite imagery and social media images Alam et al. (2018). Model performance is typically evaluated using metrics such as accuracy and F1 score.
Estimation datasets usually provide quantitative labels such as flood depths Akinboyewa et al. (2024). Metrics like Mean Absolute Error (MAE) are used for evaluation.
Generation datasets are also extensively used and primarily fall into two categories: question answering and summarization. Question-answering datasets provide disaster-related questions paired with crowdsourced annotated answers Rawat (2024). Additionally, multimodal question-answering datasets, which incorporate disaster-related images as contextual information, are widely utilized Sun et al. (2023). For the summarization task, large collections of documents serve as inputs, with reference summaries curated by domain experts McCreadie and Buntain (2023). Both question-answering and summarization tasks are evaluated using metrics such as BLEU.
Extraction datasets identify and label specific elements within a sentence, such as keywords Nguyen and Rudra (2022a) and locations Suwaileh et al. (2022). Tokens are labeled as "outside," "start," or "end" to indicate their extraction status. These datasets are primarily used for token-level classification tasks and are evaluated using classification metrics.